the length of a rectangle is 3times its worth. if the perimeter is 72cm. calculate the width of the rectangle
Answer by macston(5194)

W=width; L=length=3W P=perimeter=2(L+W)=72 cm
P=2(L+W) Substitute for L
72cm=2(3W+W)
72cm=8W divide each side by 8
9cm=W ANSWER The width of the rectangle is 9 cm.
CHECK:
L=3W=3(9cm)=27cm
P=2(L+W)
72cm=2(27cm+9cm)
72cm=2(36cm)
72cm=72cm